OpenStudy (eva12):

1)integral xcos(1x+y) dx help please

OpenStudy (anonymous):

i dunno. but i love ur prophile photo!

OpenStudy (anonymous):

check your statement.

OpenStudy (nincompoop):

\[\int\limits x \cos (1^x+y) dx? \]

OpenStudy (eva12):

\[\int\limits_{0}^{\pi/4}\int\limits_{0}^{\pi/6}xcos(1x+y) dxdy\] its a double integral

ganeshie8 (ganeshie8):

switch it to dydx

OpenStudy (nincompoop):

isn't 1x just x?

ganeshie8 (ganeshie8):

\(\int\limits_{0}^{\pi/4}\int\limits_{0}^{\pi/6}xcos(1x+y) dxdy = \int\limits_{0}^{\pi/6} \left(\int\limits_{0}^{\pi/4} xcos(1x+y) dy\right )dx\)

ganeshie8 (ganeshie8):

lol yes, i think the integrand has to be "xcos(x+y)" "xcos(1x+y)" makes me feel it has a typo or something..

OpenStudy (eva12):

nope there is no typo its a webwork question

ganeshie8 (ganeshie8):

okay, reverse the order and evaluate ?

OpenStudy (shiraz14):

Use integration by parts - sub. u=x; dv=cos(x+y)dy.